# Welfare Analysis in Dynamic Models

### v4

@@ 1. Introduction
The economic motivation for the welfare decomposition we propose is drawn from various fields in applied economics, including labor economics, public finance, and health economics. In labor economics, Kitagawa (1955), Oaxaca (1973) and Blinder (1973) decompose the gender wage gap into structural and composition effects. In public finance, Chetty (2009) distinguishes between mechanical and behavioral  effects of taxation policies. In health economics, the work by Einav et al. (2010)  differentiates the direct effect of a health insurance price change from the indirect effect arising from changes in selection. For an example of cancer screening (Einav et al., 2020), the total welfare effect corresponds to the introduction of a new screening option, and the direct effect is its counterpart as if the frequency of screening were mandated. One of the paper’s contribution is to give identification, estimation, and inferential results for direct and indirect effects in dynamic models.
@@
As an empirical application, we revisit the study of teacher absenteeism in the nonformal education centers (NFEs) of Rajasthan, India, originally analyzed by Duflo et al. (2012) (henceforth, DHR). First, we replicate their structural estimates using an infinite-horizon dynamic binary choice model as in Aguirregabiria and Mira (2002). Our estimates match those of DHR, which solves the finite-horizon model,  up to the third decimal place. This exercise demonstrates that our methodological results have broader empirical relevance: when the panel is sufficiently long, the distribution of the state vector can be treated as approximately stationary in the middle of the panel. Second, we incorporate  prior work history as a determinant of leisure utility. While these factors has strong predictive power, the structural estimates remain close to those reported by DHR. Finally, we estimate average welfare across different teacher groups determined by their score before they begin teaching.  Our welfare estimates exhibit interpretable patterns and are robust to the choice of regularized estimators.
@@ 2. Setup
Let $\mathcal{X}$ be a set and $X$. Define
@@ 3. Decomposition of Differences in Average Welfare
To motivate our analysis, we present a simple running example in the context of a randomized controlled trial, in which a one-shot treatment assigned at time $t=0$ generates dynamic incentives. Suppose we aim to analyze welfare differences between the treated and control populations, denoted by $1$ and 00, respectively. In each group, the time-varying state variable is denoted by $S$, and the full state vector is $X=(S,K)$.
@@
reflects changes in agents’ behavior that alter the distribution of states, holding the utilities. This term can be interpreted as the indirect or behavioral effect. The sum of the direct and behavioral effects gives the total welfare difference.
@@
The study by Duflo et al. (2012) estimates a dynamic behavioral structural model in which teachers choose between working and taking leisure, with the goal of predicting their choices as a function of cash incentives for attendance. In contrast to their focus, we examine teacher factual and counterfactual welfare.
Let $K=1$ indicate assignment to a cash bonus for daily attendance. The state variable $S$ denotes the number of days attended since the beginning of the month, and the full state vector is $X=(S,K)$. The direct effect, $\delta_{\langle 1\mid 1\rangle}-\delta_{\langle 1\mid 0\rangle}$, captures the impact of the bonus on welfare through per-period utilities, holding attendance behavior fixed. The indirect effect, $\delta_{\langle 1\mid 0\rangle}-\delta_{\langle 0\mid 0\rangle}$, reflects changes in stationary attendance patterns (i.e., choice probabilities) induced by the treatment, holding per-period utilities fixed. A similar decomposition applies to the dynamic behavioral model studied in Todd and Wolpin (2006).
@@
A standard approach to evaluating welfare in the context of cancer screening focuses on mortality outcomes. However, a broader perspective considers the costs of screening—such as time, resource use, and the psychological burden of false positives—which are naturally captured by welfare metrics based on latent utilities inferred from observed choices.
@@
Proposition 3.1 expresses the counterfactual welfare $\delta_{\langle 1\mid 0\rangle}$ as a linear functional of the treated value function that does not depend on the stationary distributions.
@@
Recent work has extended classical decomposition methods to modern settings. Chernozhukov et al. (2022) derive the Riesz representer for the Average Treatment Effect on the Treated (ATET), and Vafa et al. (2024) provides an Oaxaca-Blinder decomposition of wage differences. Proposition 3.1 departs from these approaches by offering a decomposition of average welfare in dynamic models where welfare is based on latent utilities rather than observed outcomes.
@@
We study how daily financial incentives affect teacher attendance in single-teacher nonformal education centers (NFEs) operated by the NGO Seva Mandir in tribal villages of Udaipur, Rajasthan, India. From 2003 to 2005, Duflo et al. (2012) conducted a randomized trial in which tamper-proof cameras recorded photos at school opening and closing. A school day was deemed valid if the two images were at least five hours apart and showed at least eight students present. At the end of each month, teachers earned a base salary of 500 Rupees (Rs) per month if they worked fewer than 10 days, plus a 50 Rs bonus for each additional day of work. The 10-day threshold for bonus activation created the dynamic incentive, which we focus on in this application [^2] We abstract from the firing threat, which appears negligible: no teacher was fired during the study period, even in cases of near-total absence. According to Duflo et al. (2012), Seva Mandir takes a long-term view in assessing teacher performance, which may explain the lack of dismissals..  Our dataset consists of daily attendance records for 57 teachers over an 18-month period, from January 1, 2004, to June 30, 2005, and a test score obtained from an exam before teaching. [^3] Following Duflo et al. (2012), the dataset includes only days when teachers actively choose between work and leisure. Holidays and weekends, though automatically counted toward teachers’ total days worked, are excluded from the estimation sample.
4.1. Replicating Duflo et al. (2012)
@@
Here, $\bar{x}_{t}\in\mathrm{R}^{p_{X}}$ is the state vector including a constant and, depending on the model, may include other observable teacher characteristics, and $\mu_{0}\in\mathrm{R}^{p_{X}}$ is a parameter to be estimated. The state vector is $X_{t}=(\bar{X}_{t},d_{t})$.   For example, in Model I of DHR, the leisure utility is assumed to be the same for all teachers, which corresponds to $u(x_{t},0)=\mu_{0}$ and $X_{t}=d_{t}$.  The per-period utilities are not discounted. The number of components in $\bar{x}_{t}$ is taken to be sufficiently small so that maximum likelihood estimators could be applied.
We consider a stylized dynamic binary choice model that is a special case of the one described in Section 2. Since teachers cannot be fired, the base salary should not affect their choice to work.  We decompose the total monthly bonus into daily payments.  In other words,  we consider the model where the utility of working ($j_{t}=1$) on day $t$ is given by
@@
where the indicator function, referred to as “In the money” in DHR,  reflects the daily structure of the bonus in the stylized model.   The parameter $\mu_{1}$ converts consumption, measured in Rupees, into utility terms. The utility of leisure is the same as in equation (4.2). The stylized model (4.3)–(4.2) has the same monetary flow as the exact model (4.1)–(4.2),  up to discounting of the bonus. To abstract away from the finite horizon effects, we restrict the sample to calendar days 15, 16, and 17 of each month. In the absence of finite horizon restriction, the methods proposed in this paper accommodate a rich pattern of observable heterogeneity in (4.2) that would have been prohibitive for likelihood estimators.
Table 1  compares  exact DHR estimates (Columns (1)-(2))  to their infinite-horizon replicates (Columns (3)-(4)) for selected coefficients. First, the estimated bonus effect $0.039$ lies within the range 0.0160.0160.0160.0160.0490.0490.0490.049.  On average, the replicated standard errors exceed the exact ones by a factor ranging from $2.5$ to $3$.  Since only three days from each month are used, the drop in the precision is expected. Second, the sign of the teacher test score coefficient remains consistent and negative across specifications. This preliminary evidence suggests that the finite horizon effects are negligible, and the methods developed in this paper are applicable to this dataset.
4.2. Adding State Variables.
The utility of skipping work is difficult to quantify in Rupees. As mentioned by DHR, the choice to take leisure may depend on social norms and informal requests, past effort, or fatigue. The observables such as the teachers test score may not adequately capture teacher heterogeneity. Following DHR’s suggestion, [^4] On page 1225, footnote 16 Duflo et al. (2012) mention that teachers decision to work may be influenced by prior work history. we include a vector of prior work history in $\bar{x}_{t}$ as a state component. We anticipate that the work history is sufficiently rich to adequately represent teachers preference for leisure. In addition, we summarize this vector using a work streak variable defined as the number of consecutive days a teacher has worked without taking leisure. Formally, the work streak is
@@
Table 2 shows the out-of-sample mean squared error (MSEs) for predicting a teacher’s decision to work, using models that sequentially expand the covariate set from Model I to Model V. Adding the teacher’s test score (Model II) yields minimal improvement. Including month dummies (Model III) results in a modest reduction in MSE. Incorporating the work streak (Model IV) leads to a substantial improvement: relative to Model III, the MSE falls by roughly 32% for Logit, 32% for Probit, and 35% for Random Forest. Finally, adding the full 90-day work history dummies (Model V) further improves prediction, halving the MSE relative to the baseline. These findings are consistent with DHR’s hypothesis that prior work history is relevant for capturing teacher’s decision to work, even though they do not reveal which aspect of the per-period utilities prior work history is relevant for.
Table 1. Results from the Exact (DHR) and Stylized Structural Models
@@
Teacher Test Score $(\mu_{0,Score})$0.005 -0.011
@@
Notes: Table reports selected structural estimates for the dynamic behavioral model. Columns (1)-(2) replicate DHR. Columns (3)–(4) estimate a stylized model defined in equations (4.3)–(4.2). For the latter case, the sample consists of 57 teachers observed over 18 months as well observations from the middle days of each month (15, 16, and 17). The discount factor is $\beta=0.99$. See main text for details.
@@
I Days Worked 0.174 0.175 0.153
II $+$ Test Score 0.173 0.174 0.153
III $+$ Month Dummies 0.171 0.171 0.147
IV $+$ Work Streak 0.116 0.116 0.095
V $+$ Work History Dummies 0.062 0.063 0.069
Notes: This table reports the Mean Squared Error (MSE) for predicting teacher decision to work, using models that progressively add explanatory variables. Row 1 includes only the number of days worked in the current month. Row 2 adds the teacher’s test score. Row 3 introduces month dummies. Row 4 includes the work streak, defined as the number of consecutive days worked since the teacher’s last absence. Row 5 adds the full 90-day work history vector. Models are trained on 2004 data and evaluated on 2005 observations. Columns correspond to Logistic Regression (Logit), Probit Regression (Probit), and Random Forest (RF). See Table 4 for descriptive statistics of the variables.
4.3. Teacher Welfare.
We model the utility of leisure as a linear function
@@
where $\bar{x}_{t}$ includes a constant, the teacher’s test score, and a vector of 90-day work history dummies. This specification allows for heterogeneity in leisure preferences based on recent work history. The structural parameters $(\mu_{1},\mu_{0})$ include 92 coefficients, which we estimate using Algorithm 1 in Appendix G. This algorithm requires conditional choice probabilities and several nuisance components, including parts of the value function and the dynamic dual representation. For choice probabilities, we consider both logistic regression and random forest estimators, which correspond to the columns of Table 3. The value function components are estimated via the Lasso-based method described in Section 7.
Table 3 presents our main empirical findings. Panel A reports the structural estimates, while Panels B through D present average welfare results. Columns (1)–(2) correspond to a simple leisure utility specification where $\bar{x}_{t}$ includes only a constant and test score (Model II in Table 2). Columns (3)–(4) correspond to the full specification that includes work history dummies (Model V). Structural estimates are similar across specifications and remain close to those in Duflo et al. (2012). As expected, standard errors increase in the richer specification due to the higher number of coefficients being estimated.
The primary objective of Duflo et al. (2012) is to assess the effect on student outcomes. Yet, it is equally important to evaluate the welfare of teachers whose professional lives were affecte by the experiment. According to Duflo et al. (2012), several teachers reported improved discipline and focus due to monitoring and cash incentive. However, their original study did not provide quantitative welfare measures. We fill this gap by reporting average welfare in Panels B, C, and D of Table 3.
To estimate welfare, we use the dual estimator described in Section 6. Instead of estimating value function, the estimator combines the structural parameters from Panel A and the estimated choice probabilities. Across all models, the average welfare ranges from 165 to 188 Rupees per period. These patterns are consistent with observed attendance rates (Table 4) and the negative sign of the test score coefficient in Panel A. In particular, lower-scored teachers receive consistently lower welfare estimates than their higher-scored counterparts. This suggests that policymakers may wish to consider setting a minimum test score threshold, making the bonus conditional on skill, or both.
@@
Panel A: Estimates for Selected Coefficients in the Structural Model
“In the money” ($\mu_{1}$) 0.039 0.039 0.040 0.040
@@
Teacher Test Score $(\mu_{0,Score})$0.011 -0.011 -0.017 -0.017
@@
Notes: Table reports the welfare metrics for the dynamic behavioral model defined in equations (4.3)–(4.2). Panel A reports structural estimates. The parameter in Panel B (Average Welfare) is defined in Example 2.1. The parameters in Panels C and D (Group Average Welfare) are defined in Example 2.2. Models II and V correspond to the state components in Table 2. Teachers are classified as Low-Scored if their test score is less than or equal to 30, and High-Scored if their score is greater than or equal to 40. The discount factor is $\beta=0.99$. See main text for further details.

### v5

@@ 1. Introduction
The economic motivation for the welfare decomposition we propose is drawn from various fields in applied economics, including labor economics, public finance, and health economics. In labor economics, Kitagawa (1955), Oaxaca (1973) and Blinder (1973) decompose the gender wage gap into structural and composition effects. In public finance, it is common to distinguish between mechanical and behavioral  effects of taxation policies, see e.g. Chetty (2009). In health economics, the work by Einav et al. (2010)  differentiates the direct effect of a health insurance price change from the indirect effect arising from changes in selection. For an example of cancer screening (Einav et al., 2020), the total welfare effect corresponds to the introduction of a new screening option, and the direct effect is its counterpart as if the frequency of screening were mandated. One of the paper’s contribution is to give identification, estimation, and inferential results for direct and indirect effects in dynamic models.
@@
We revisit the study of teacher absenteeism in the nonformal education centers (NFEs) of Rajasthan, India, initially analyzed by Duflo et al. (2012) (henceforth, DHR). Their approach emphasizes unobserved heterogeneity in teachers’ leisure preferences—shaped by factors such as past effort, illness, fatigue, and informal obligations. We conjecture that this heterogeneity can be effectively proxied by a sufficiently long window of prior work history, even if the history itself may not have a structural or causal interpretation. DHR’s structural estimates remain robust to this exercise. In addition, we focus on teacher welfare and find that the failure to properly account for teacher heterogeneity results in the average teacher welfare overestimated by 13–20%%. As a side contribution, we demonstrate that our methodological framework remains valid in finite-horizon settings, provided the panel is sufficiently long relative to standard discount factors.
@@ 2. Setup
Let $\mathcal{X}$ be a set and $X$ be a state vector. Define
@@ 3. Decomposition of Differences in Average Welfare
To motivate our analysis, we present a simple running example in the context of a randomized controlled trial, in which a one-shot treatment assigned at time $t=0$ generates dynamic incentives. Suppose we aim to analyze welfare differences between the treated and control populations, denoted by $1$ and 00, respectively. In each group, the time-varying state variable is denoted by $S$. The full state vector is $X=(S,K)$ where $K$ is the indicator of the treatment status. We adopt the notation in Chernozhukov et al. (2013).
@@
reflects changes in agents’ behavior that alter the distribution of states, holding the utilities fixed. This term can be interpreted as the indirect or behavioral effect. The sum of the direct and behavioral effects gives the total welfare difference.
@@
Duflo et al. (2012) estimates a dynamic behavioral structural model in which teachers choose between working and taking leisure. Duflo et al. (2012) focuses the student achievements as the primary target. In this paper, we take a complementary perspective and focus on teacher welfare.
Let $K=1$ indicate the treatment status, where treatment constitutes a cash bonus for each additional day of work once the count of days worked exceeds 10 in a given month. The state variable $S$ denotes the number of days worked since the beginning of the month. The full state vector is $X=(S,K)$. The direct effect, $\delta_{\langle 1\mid 1\rangle}-\delta_{\langle 1\mid 0\rangle}$, captures the impact of the bonus on welfare through per-period utilities, holding teachers work decisions fixed. The indirect effect, $\delta_{\langle 1\mid 0\rangle}-\delta_{\langle 0\mid 0\rangle}$, arises from the change in teachers work decisions, holding per-period utilities fixed. A similar decomposition applies to the model in Todd and Wolpin (2006).
@@
A standard approach to evaluating welfare in the context of cancer screening focuses on mortality. Yet, a broader perspective considers the costs of screening—such as time, resource use, and the psychological burden of false positives— which cannot be captured by observable outcomes.
@@
Proposition 3.1 expresses the counterfactual welfare $\delta_{\langle 1\mid 0\rangle}$ as a linear functional of the treated value function.
@@
Recent work has extended classical decomposition methods to modern settings. Chernozhukov et al. (2022) derive the Riesz representer for the Average Treatment Effect on the Treated (ATET). Vafa et al. (2024) provides an Oaxaca-Blinder decomposition of wage differences. Proposition 3.1 departs from these approaches by offering a decomposition of average welfare in dynamic models where welfare is based on latent utilities rather than observed outcomes.
@@
We study how daily financial incentives affect teacher attendance in single-teacher nonformal education centers (NFEs) operated by the NGO Seva Mandir in tribal villages of Udaipur, Rajasthan, India. From 2003 to 2005, Duflo et al. (2012) conducted a randomized trial in which tamper-proof cameras recorded photographs at school opening and closing. A school day was deemed valid if the two images were at least five hours apart and at least eight students were present. At the end of each month, teachers earned a base salary of 500 Rupees (Rs) if they worked fewer than 10 days, plus a 50 Rs bonus for each additional day of work beyond that threshold. The 10-day cutoff thus created a nonlinear dynamic incentive, which we focus on in this application. [^2] We abstract from the firing threat, which appears negligible: no teacher was fired during the study period, even in cases of near-total absence. According to Duflo et al. (2012), Seva Mandir adopts a long-term view in assessing teacher performance, which may explain the lack of dismissals. Our dataset comprises daily attendance records for 57 teachers over an 18-month period (January 1, 2004 to June 30, 2005), along with a test score administered prior to the start of teaching. [^3] Following Duflo et al. (2012), the estimation sample includes only weekdays when teachers actively choose between working and taking leisure. Holidays and weekends, though counted toward pay, are excluded from analysis.
@@
Here, $\bar{x}_{t}\in\mathrm{R}^{p_{X}}$ is the state vector including a constant and possibly other observable characteristics, and $\mu_{0}\in\mathrm{R}^{p_{X}}$ is a parameter to be estimated. The state vector is $X_{t}=(\bar{X}_{t},d_{t})$.   For example, in Model I of DHR, the leisure utility is assumed to be the same for all teachers, which corresponds to $u(x_{t},0)=\mu_{0}$ and $X_{t}=d_{t}$.  The per-period utilities are not discounted. DHR includes only a handful of observables into $\bar{x}_{t}$ so as to leverage standard maximum likelihood estimators.
We consider a stylized dynamic binary choice model as described in Section 2. Since teachers cannot be fired, the base salary is assumed not to affect their choice between work and leisure.  We decompose the total monthly bonus into daily payments.  Specifically, we assume the utility of working ($j_{t}=1$) is given by:
@@
where the indicator function, referred to as “In the money” by Duflo et al. (2012), captures the bonus structure in the stylized model.   The parameter $\mu_{1}$ converts monetary rewards (in Rupees) into utility units. The stylized model in equations (4.3)–(4.2) preserves the monetary incentives of the exact model (4.1)–(4.2), aside from discounting of the bonus. To abstract from the finite-horizon considerations, we restrict attention to calendar days 15, 16, and 17 of each month. If this model is relevant, the methods developed in this paper permit the state vector $\bar{X}_{t}$ to include high-dimensional covariates.
Table 1 compares the exact estimates reported by Duflo et al. (2012) (Columns (1)-(2)) to our stylized infinite-horizon replications (Columns (3)-(4)) for selected coefficients. The results are encouraging. First, the estimated bonus coefficient under the stylized model falls within the range of the exact estimates. The standard errors of the replicated coefficients are, on average, 2.5 to 3 times larger, as expected given the smaller sample (only three days per month). Second, the coefficient on teacher test scores remains negative across all specifications, consistent with the original findings. Other coefficients (not reported here) also closely match their exact counterparts. These results suggest that the stylized infinite-horizon framework is appropriate for this dataset.
We further investigate the role of prior work history in explaining teachers work decisions, a possibility raised by Duflo et al. (2012, footnote 16). Duflo et al. (2012) included the first lag of work history in Model VIII, and we extend this idea by incorporating 89 more lags. Additionally, we summarize work history using a work streak variable, defined as the number of consecutive days a teacher has worked without taking leisure:
@@
We investigate the role of prior work history both in the predictive and structural settings.
Table 2 shows the out-of-sample mean squared error (MSEs) for predicting a teacher’s decision to work, using models that sequentially expand the covariate set from Model I to Model IV. Adding the teacher’s test score on top of days worked yields minimal improvement. Including month dummies (Model II) results in a modest reduction in MSE. Incorporating the work streak (Model III) leads to a substantial improvement: relative to Model II, the MSE falls by roughly 32% for Logit, 32% for Probit, and 35% for Random Forest. Finally, adding the full 90-day work history dummies (Model IV) further improves prediction, halving the MSE relative to the baseline. These findings show that prior work history has high predictive power even after other observables have been taken into account.
As a next exercise, we revisit the stylized structural model with the aim of flexibly modeling the utility of leisure. As noted by Duflo et al. (2012), decisions to skip work may be influenced by factors such as social norms, informal requests and commitments, accumulated effort, or fatigue. These factors are unlikely to be fully captured by basic observables like test scores. A longer spell of prior work history may be more granular and thus may better represent teacher’s heterogeneity even if the history itself has no structural or causal interpretation. Assuming only a small number of lags suffices to capture the history, we include a 90-day window of lagged attendance indicators. This sparsity assumption calls for the use of Lasso estimators of the value function and the dynamic dual representation, developed in Section 7.
Table 1. Structural Estimates from Exact (DHR) and Stylized (CNS) Models
@@
Teacher Test Score -0.005 -0.011
@@
Notes: Table reports selected coefficients for structural estimates. Columns (1)-(2) replicates exact, finite-horizon, estimates from Models I and VIII using the full sample of 57 teachers observed over 18 months where payoffs are not discounted. Columns (3)–(4) report their analogs for the stylized model defined in equations (4.3)–(4.2) with distinct choices of state components. The stylized, infinite-horizon, model is estimated using only the observations from the middle days of each month (15, 16, and 17) and the discount factor is $\beta=0.99$. See main text for details.
@@
I Days Worked $+$ Test Score 0.173 0.174 0.153
II $+$ Month Dummies 0.171 0.171 0.147
III $+$ Work Streak 0.116 0.116 0.095
IV $+$ Work History Dummies 0.062 0.063 0.069
Notes: Table reports the Mean Squared Error (MSE) for predicting teacher decision to work, using models that progressively add explanatory variables. Row I includes only the number of days worked in the current month and the teachers test score, Row II adds month dummies, and so on. Models are trained on 2004 data and evaluated on 2005 observations. Columns correspond to Logistic Regression (Logit), Probit Regression (Probit), and Random Forest (RF). See Table 4 in Appendix for descriptive statistics of the variables.
Table 3 reports selected coefficients for the structural parameter estimates (Panel A) as well as welfare metrics (Panels B, C and D). Columns (1)-(2) correspond to a simple model of (4.2) whose only observable is teacher’s test score. Columns (3)-(4) correspond to a more sophisticated model of (4.2) where observables include prior work history. In both cases, the model is estimated using Algorithm 2 described in Appendix G based on Logit and Random Forest estimators of conditional choice probability. The welfare metrics are estimated using the dual estimator described in Section 6. Instead of using the value function, it combines the structural estimates of Panel A with the choice probabilities.
@@
Panel A: Structural Estimates
Bonus ($\mu_{1}$) 0.039 0.039 0.049 0.046
@@
Teacher Test Score -0.011 -0.011 -0.012 -0.013
@@
Notes: Table reports findings for the stylized model defined in equations (4.3)–(4.2). The choice of components corresponds to Models I and IV in Table 2. Panel A reports selected coefficients of structural parameter estimated using Algorithm 2 in the Appendix. Panels B, C and D report welfare metrics defined in Examples 2.1 and 2.2. Teachers are classified as Low-Scored if their test score is less than or equal to 30, and High-Scored if their score is greater than or equal to 40. See main text for further details.